OpenStudy (anonymous):

HELP! PLEASE! ASAP a triangle has angles D, E, and F. Which of the following could not be a set of angles? A. m

OpenStudy (anonymous):

I think It's D

OpenStudy (anonymous):

no it's A because 45+45=90 and sum of 2 angles of a triangle should be greater than the 3rd one

OpenStudy (anonymous):

because \(\Large \angle D+\angle E +\angle F=90+46+46=182 (>180) \)

OpenStudy (anonymous):

So the answer is D?

OpenStudy (anonymous):

@dhirensharma you should look at the question

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Thank you! (:

OpenStudy (anonymous):

it's A fr god's sake jung that is the basic rule fr a triangle to be a triangle the msum of the measures of any of it's two sides should be greater than the measure of the 3rd side are u kidding me tthat's the basic rule 45+45=90 but it aint greater than 90 therefore it's AAAAA

OpenStudy (anonymous):

srry it's d i get it

OpenStudy (anonymous):

"could not be a set of angles"

OpenStudy (anonymous):

46+46+90 is more than 180 thats wy so it's d D